OpenStudy (anonymous):

An archeologist discovered two artifacts buried beneath the ground. She found the first artifact at an elevation of -8 2/7 meters and the second at an elevation of -35 4/7 meters. What is the distance between these two elevations? -43 6/7 meters -27 2/7 meters 27 2/7 meters 43 6/7 meters

OpenStudy (anonymous):

@kaylie.sullivan can you help me with this problem please

OpenStudy (danjs):

They want the difference (subtraction) between those two mixed numbers

OpenStudy (anonymous):

how?? -8 2/7 - (-35 4/7)= ???

OpenStudy (tinylittlehelper):

Hint: Distance is also called absolute value. And yes, that is correct.

OpenStudy (danjs):

mixed numbers are strange to work with i think, they are 2 numbers added together, maybe put them into improper fractions first, they are both already over the same denominator

OpenStudy (tinylittlehelper):

But: -a - (-b) = -a + b

OpenStudy (anonymous):

now I understand!!! so the answer is 27 2/7???

OpenStudy (danjs):



OpenStudy (tapha):

c man..

OpenStudy (danjs):

the distance is a positive measurement

OpenStudy (anonymous):

Thanks!!!!!

OpenStudy (danjs):

\[\sqrt{(a-b)^2}\] a-b is a squared quantity, always positive

OpenStudy (danjs):

welcome

OpenStudy (tinylittlehelper):

\[\sqrt{(a-b) ^{2}} = |a-b|\]

OpenStudy (danjs):

yes